Solve the differential equation of the following linear differential equations. Show complete solution.


dy/dx + y = e^2x


1
Expert's answer
2021-09-30T02:03:33-0400

The integration factor: "\mu(x)=e^x"


"e^x\dfrac{dy}{dx}+e^xy=e^{3x}"

"d(e^xy)=e^{3x}dx"

Integrate


"e^xy=\int e^{3x}dx"

"e^xy=\dfrac{1}{3} e^{3x}+C"

"y=\dfrac{1}{3}e^{2x}+Ce^{-x}"